A screw jack has a single start thread of pitch [p] 6 mm and an operating handle of effective radius [r] 450 mm. It is found that an input effort [E] of 45 N at the handle is required to raise a load [W] of 8 kN. For the above conditions, Determine :
(i) The velocity ratio
(ii) The mechanical advantage
(iii) The operating efficiency
(iv) Also give a reason, whether the jack is likely to overhaul or not.
1
Expert's answer
2018-04-19T09:45:08-0400
(i) The velocity ratio
VR=2πr/p=2π 450/6= 471.24
(ii) The mechanical advantage
MA=W/E=8000/45=177.78
(iii) The operating efficiency
η=MA/VR=177.78/471.24=0.377 or 37.7%
(iv) A machine will overhaul if its efficiency is greater than 50%. Thus, the jack is not likely to overhaul.